OpenStudy (anonymous):

Will Medal!!!! What is the point-slope form of the equation of a line that passes through the point (3, 5) and has a slope of 2? A.y+3=2(x+5) B.y-3=2(x-5) C.y-5=2(x-3) D.y-5=-2(x-3)

OpenStudy (crashonce):

remember the slope form is y=mx+b

OpenStudy (crashonce):

or use y=y1=m(x-x1)

OpenStudy (crashonce):

so ur equation is y-5=2(x-3)

OpenStudy (crashonce):

C

OpenStudy (anonymous):

Thanks!!! @CrashOnce